OpenStudy (anonymous):

Wendy is comparing two checking accounts. One has a monthly fee of $9 and a per-check fee of $0.50, and the other has a monthly fee of $5 and a per-check fee of $0.55. What is the minimum number of checks Wendy needs to write for the first bank to be a better option?

OpenStudy (anonymous):

9 + .50x = 5 + .55x .50x - .55x = 5 - 9 -.05x = - 4 x = -4/-.05 x = 80 80 checks would have to be written for them to be EQUAL...therefore, for the first bank to be a better option, there has to be 81 checks written. check... (A) 9 + .50(81) = 9 + 40.50 = 49.50 <-- better option (B) 5 + .55(81) = 5 + 44.55 = 49.55